a man who with 55kg, on a skateboard. He had an

initial speed of 2.77m/s when descending the stairs at a height of 2m.

Calculate the man’s potential energy when his speed is 6m/s


1
Expert's answer
2021-01-22T09:56:41-0500
"E_p+0.5mv^2=mgh+0.5mv^2_0\\
E_p=mgh+0.5m(v^2_0-v^2)\\E_p=(55)(9.8)(2)+0.5(55)(2.77^2-6^2)=299\ J"